OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. a = ?, b = 13, c = 21

OpenStudy (anonymous):

You can use the pythagorean thereom the a^2+b^2=c^2 Just plug in the numbers and solve for a a^2+13^2=21^2

OpenStudy (wolf1728):

Since a² + b² = c² then a² = c² -b² a² = 21² - 13² a² = 441 - 169 I just thought I'd add my post to make the solution a little easier to follow. Alex108 - I think you can finish the problem on your own.

OpenStudy (anonymous):

a squared plus b squared equals c squared therefore a squared equals c squared - b squared 21 squared - 13 squared equals 272 then square root 272 = a a = 16.49